1. An iron plate 2cm thick has a cross sectional of 5000cm^2. One face is at 150°C and the other is at 140°C. How much heat passes through the plate each second? The thermal conductivity for iron is kT= 80 W/m. K
1
Expert's answer
2018-08-27T09:43:09-0400
The Fourier's law states
Q/Δt=k_T A ΔT/d
So
Q/Δt=80×5000×〖10〗^(-4)×(150-140)/(2×〖10〗^(-2) )=20000 W=20 kW